it works out the demand function for the book as:                                              

Q = 5000 – 5P

Find out

i)                   Demand curve

ii)                 Number of book sold at P = Rs. 25

iii)               Price for selling 2500 copies

iv)               Price for zero sales

v)                 Elasticity for fall in price from Rs. 25 to Rs. 20.


1
Expert's answer
2021-04-27T07:43:15-0400

i)

In the demand function, Price (P) is expressed in terms of Quantity (Q).

Since "Q = 5000 - 5P"

"5P = 5000 - Q"

"P = 1000 - 0.2Q"

"P = 1000 - 0.2Q"   is the demand curve


ii)

"P=1000-0.2Q"

"25=1000-0.Q"

"0.2Q=1000-25"

"0.2Q=975"

"Q=4875"


iii)

Q=2500

"P=1000-0.2Q"

"P=1000-0.2(2500)"

"P=1000-500"

"P=500"


iv)

Q=0

"P=1000-0.2Q"

"1000-0.2(0)"

P=1000


v)

P=25 Q=4875

P=20

"20=1000-0.2Q"

"0.2Q=1000-20"

"0.2Q=980"

"Q=4900"

P25 Q 4875

P20 Q 4900

Price elasticity of demand"=\frac {(Q_2-Q_1)/[(Q_2+Q1)/2]}{(P_2-P_1)/[(P_2+P_1)/2}"


"=\frac{(4900-4845)/[(4900+4875)/2]}{(20-25)/[(20+25)/2]}"

"=\frac{0.0051}{0.22}"

"=0.023"